OpenStudy (anonymous):

at a run festival 3/4 of the runners are woman. if there is 560 runners, how many are runners. Type a whole number.. can some help me with this problem.

OpenStudy (anonymous):

can someone help me with the problem above i wrote

OpenStudy (anonymous):

is ur question correct coz u have already mentioned number of runners and asking for the same.

OpenStudy (anonymous):

Not sure what you are asking, retype your question

OpenStudy (anonymous):

total number of runners = 560 According to the questrion, number of women = \[\frac{ 3 }{ 4 } \times (Total Runners)\]\[\frac{ 3 }{ 4 } \times 560 \] = 420 And the number of men = 560 -420 = 140.